OpenStudy (anonymous):

Use an appropriate local linear approximation to estimate the value of the given quantity. (3.02)^4 how do i do this

myininaya (myininaya):

We could use the function f(x)=(x+3)^4 to help an approximating (3.02)^4 or even f(x)=x^4

myininaya (myininaya):

So we know the tangent line should look like \[y-f(a)=f'(a)(x-a) \] So say we used \[f(x)=(x+3)^4 \] And say we wanted to find the tangent line at a=0

OpenStudy (anonymous):

how do you get that

myininaya (myininaya):

So we would have \[y=f'(a)(x-a)+f(a) \\ y=f'(0)(x-0)+f(0) \\ f(x) =(x+3)^4 \approx f'(0)(x-0)+f(0) \text{ for values near x=0 } \]

myininaya (myininaya):

How do you get what?

OpenStudy (anonymous):

what happens to .02

OpenStudy (anonymous):

and why the zero?

myininaya (myininaya):

We will approximate (0.2+3)^4 in a sec

myininaya (myininaya):

I chose 0 because 0.2 was close to the integer 0

myininaya (myininaya):

oops I mean 0.02 every where I had 0.2

myininaya (myininaya):

Let me do one example all the way though. Say we wanted to do a linear approximation for sqrt(10.1) There are different functions you could choose to help you do this. But I definitely want my function to involve a sqrt Say I want to find the tangent line to the curve \[f(x)=\sqrt{x+10} \] at x=0... \[y=f(0)+f'(0)(x-0) \\ y= \sqrt{10}+\frac{1}{2 \sqrt{x+10}} |_{x=0} (x-0) \\ y=\sqrt{10} +\frac{1}{2 \sqrt{10}} (x-0) \\ y=\sqrt{10}+\frac{x}{2 \sqrt{10}} \\ \sqrt{x+10} \approx \sqrt{10} +\frac{x}{2 \sqrt{10}} \text{ for values of } x \text{ near x=0 } \] So we wanted to approximate sqrt(10.1) so we need x=.1 (this is a value near x=0 so it should be a sorta good approximation) \[\sqrt{.1+10} \approx \sqrt{10} +\frac{.1}{2 \sqrt{10}} =\sqrt{10}+\frac{1}{20\sqrt{10}}=\frac{20 (10)+1}{20 \sqrt{10}}=\frac{201}{20\sqrt{10}}\]

myininaya (myininaya):

http://www.wolframalpha.com/input/?i=sqrt%2810.1%29 sqrt(10.1) is approx 3.178 according to a calculator and our approximation 201/[20*sqrt(10)] is http://www.wolframalpha.com/input/?i=201%2F%2820*sqrt%2810%29%29

myininaya (myininaya):

As you see 201/[20*sqrt(10)] was a pretty darn good approximation in this case

OpenStudy (zarkon):

I feel your example defeats the purpose of the technique. \(\sqrt{10.1}\) is hard to calculate without a calculator and so is \(\sqrt{10}\) . \(\sqrt{9}\) though is easy to calculate without a calculator.

myininaya (myininaya):

lol. It is still close numbers.

OpenStudy (zarkon):

yes it will give a good approximation

myininaya (myininaya):

But yeah I know what you mean.

myininaya (myininaya):

We could have chose f(x)=sqrt(x+9)

myininaya (myininaya):

9 is a perfect square

OpenStudy (zarkon):

I prefer \(f(x)=\sqrt{x}\) \[f(x)=f(a)+f'(a)(x-a)\] x=10.1 and a=9 \[\sqrt{10.1}\approx\sqrt{9}+\frac{1}{2\sqrt{9}}(10.1-9)=3.18\overline{3}\]

myininaya (myininaya):

\[g(x)=\sqrt{x+9} \\ \sqrt{x+9} \approx f(0)+f'(0)(x-0) \\ \sqrt{x+9} \approx \sqrt{9} +\frac{1}{2 \sqrt{x+9}} |_{x=0} (x-0) \\ \sqrt{x+9} \approx 3 +\frac{1}{2 \sqrt{9}}(x-0) \\ \sqrt{x+9} \approx 3+\frac{1}{2(3)}x \\ \sqrt{1.1+9} \approx 3 +\frac{1}{6} (1.1) \\ \sqrt{10.1} \approx 3+\frac{1.1}{6} \\ \sqrt{10.1} \approx 3+\frac{11}{60}=\frac{180+11}{60}=\frac{191}{60}\]

myininaya (myininaya):

For some reason I love choosing the function so that I can find the tangent line at x=0

OpenStudy (zarkon):

nothing wrong with that

OpenStudy (anonymous):

ah i get it thank you your the best